OpenStudy (anonymous):

A tabletop is an isosceles right triangle. The length of the front edge of the table is 48 inches. What is the length w of each side edge? Round to the nearest tenth of an inch

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

sin B = AC / BC sin 45 = AC / 48 sqrt2 /2 = AC/48 AC = 48sqrt2 / 2 = 24sqrt2 = 24*1,41 = 33,84 AC = AB = 34 in